OpenStudy (anonymous):

A manufacturer of large enclosed cylindrical water tanks wants to know what dimensions the tanks should be in order to minimize the amount of material used. In order to find the dimensions, first consider a 10,000 litre tank (10 cubic metres). Using the formula for the volume of a cylinder, express the height of the tank (h metres) in terms of the radius (r metres).

OpenStudy (e.mccormick):

Do you know the formula for the volume of a cylinder?

OpenStudy (anonymous):

pi*h*r^2

OpenStudy (e.mccormick):

So, you have \(10=\pi hr^2\) with the given information. Right? So how would you "xpress the height of the tank (h metres) in terms of the radius (r metres)" at that point? You solve for what?

OpenStudy (anonymous):

Solve for radius? I was unsure of what is meant by "express in terms of" when related to formulas.

OpenStudy (e.mccormick):

No, solve for Height. You want a function of height that uses the radius.

OpenStudy (e.mccormick):

So with this one, it is not too hard. Since everything on the right hand side is multiplied, just divide through by everything other than height.

OpenStudy (anonymous):

Ok makes sense, cheers for that @e.mccormick

OpenStudy (e.mccormick):

The typical formula is y in terms of x, so y=x+2 or whatever. Whenever it is a in terms of b, the a part is alone and the b gets mixed with everything else. If you remember that much, that part won't be too hard. In fact, it follows the order we typically put the = sign. f(x)=x+2, where f(x) means y is... so the y side in terms of the right hand side.

OpenStudy (anonymous):

So h = 10/(pi)r^2

OpenStudy (e.mccormick):

Yah. That is how it would typically be written.

OpenStudy (anonymous):

Awesome dude, question well and truly answered.

OpenStudy (e.mccormick):

Good! Have fun!

OpenStudy (e.mccormick):

Now that I have my \(\LaTeX\) working properly again... \[\large h = \frac{10}{\pi r^2 }\]

OpenStudy (e.mccormick):

Hehe.